# Measuring AI “Slop” in Text

### v1

@@ Abstract
AI “slop” is an increasingly popular term used to describe low-quality AI-generated text, but there is currently no agreed upon definition of this term nor a means to measure its occurrence. In this work, we develop a taxonomy of “slop” through interviews with experts in NLP, writing, and philosophy, and propose a set of interpretable dimensions for its assessment in text. Through span-level annotation, we find that binary “slop” judgments are (somewhat) subjective, but such determinations nonetheless correlate with latent dimensions such as coherence and relevance. Our framework can be used to evaluate AI-generated text in both detection and binary preference tasks, potentially offering new insights into the linguistic and stylistic factors that contribute to quality judgments.
@@ 1 Introduction
“Slop” has emerged as a term describing generic, low-quality content that appears to have been generated by AI. [^1] Slop was on the shortlist of Oxford Dictionary’s Word of the Year 2024, which claims a “332% increase” in usage: https://corp.oup.com/word-of-the-year/#shortlist-2024 Recent news articles offer salient examples of such AI “slop”, ranging from non-factual claims (“… add nontoxic glue to make cheese stick to a pizza”, “geologists advise eating at least one rock a day”; Metz, 2024; Scott, 2024) to useless information (“fodder for websites whose only purpose appears to be optimising for [search engines]”; Mahdawi, 2025). Conversations on social media highlight indicators of “slop” in LLM responses, including overuse of certain terms, low information density, and structural quirks such as lists-as-responses. [^2] https://x.com/aidan_mclau/status/1884770586276381179 Despite the sudden ubiquity of the term, there is no clear definition of, nor method, for measuring “slop” in text.
@@
Our main contributions are as follows: We first introduce a working definition and taxonomy of “slop” and map each dimension to automatic metrics where possible (§3). To validate this framework, we collect span-level annotations from expert writers over 150 news articles and 100 question-answering passages to provide a fine-grained analysis of slop indicators (§4). Although binary assessments of “slop” vary across individuals, we show that our taxonomy helps explain which latent qualities (e.g., coherence, relevance, structural features) contribute most to these judgments5).
We also find that the strength of latent qualities vary based on domain, and that our taxonomy provides a useful framework for assessing quality under different tasks5). For axes that lack adequate automatic measurements (e.g., relevance, coherence, fluency), we find that standard text metrics fail to capture annotator preferences. Finally, we show that capable reasoning LLMs also fail to reliably extract and identify “slop” in text6). We provide the annotation guidelines and data at https://github.com/cshaib/slop.
@@ 2 Related Work
Text Diversity. Prior work has sought to characterize aspects of texts related to how repetitive and templated they are. Salkar et al. (2022) investigated repeated $n$-grams in LLM outputs in the context of summarization. Shaib et al. (2024b) found that modern LLMs are prone to repeatedly generate favoured syntactic templates, i.e., sequences of Part-of-Speech (PoS) tags. Padmakumar & He (2023) and Tevet & Berant (2020) examined lexical and semantic diversity in generated texts, introducing metrics to quantify variation across outputs and emphasizing its importance to generation quality. These existing efforts have informed the way in which we are thinking about what characterizes writing style and AI “slop” and provides automatic measurements for key aspects of “slop.”
Text Quality Measurements. Text quality has typically been measured using simple surface-level metrics like BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004), which can be effective when reference outputs are available (e.g., machine translation evaluation). More recent work has recognized that text quality is not monolithic but rather comprises multiple, sometimes competing dimensions that must be measured independently, and accordingly focused on multidimensional frameworks assessing properties of texts. Chakrabarty et al. (2025b) provide an editing taxonomy to correct (Chakrabarty et al., 2025a) recurring AI-writing flaws such as clichés and unnecessary exposition. Similarly, Bharadwaj et al. (2025) show that reward models over-weight 5 superficial writing cues including length, structure, jargon, sycophancy, and vagueness. Both works confirm that multiple factors contribute to text quality. Our work is complementary to measuring quality in general: We target stylistic patterns unique to LLM writing that are not covered by other taxonomies.
@@ 3 Defining LLM “Slop”
The Oxford Dictionary defines “slop” as: “[…] material produced using a large language model (LLM), which is often viewed as being low-quality or inaccurate. This type of low-quality, AI-generated material is becoming increasingly visible to people […], who often view it as unwanted or inferior.” “Slop” as a construct does not immediately permit measurement: It is difficult to quantify “low-quality” or “unwanted” text. We propose a composite measure over observable characteristics of text, where we elicit salient characteristics from a set of individuals with a range of relevant expertise. Human writing can also read as “slop”, but we adopt the above definition and focus on (seemingly) LLM-generated texts. [^3] Interestingly, what is considered “slop” in human-written text can differ in characteristics, and may even serve as an intermediary in writing processes (cf. Appendix B)
@@
Using qualitative content analysis and deductive coding techniques, we map expert definitions of “slop” to measurable concepts (Mayring, 2000). We begin by identifying key terms in survey responses and building a code list until saturation (i.e., until no new codes are created). We then map each response on to one of the following codes: Factuality, Information Density, Bias, Relevance, Repetition, Templatedness, Verbosity, Word Complexity, Tone, Coherence, Fluency, Diversity, Engagement, Vagueness, and Utility.
@@ 3.1 Datasets
We select two datasets to annotate for “slop”, motivated by two practical observations: First, that LLM-written text is becoming commonplace in reporting news on the internet (Tomlinson et al., 2025; OpenAI, 2025). The second is motivated by the use-cases outlined by our experts in §3, where a majority ($n=9$) reported using LLMs for question answering tasks (Appendix B).
News Article Generation. We evaluate “slop” over 150 news articles released by Russell et al. (2025), in which annotators are asked to label texts as being AI-written or not. Each unique article has a human-written source and a corresponding AI-written article, generated by either Claude (Anthropic, 2024), GPT-4o or o1-pro (Jaech et al., 2024). Additionally, each article includes a “humanized” article, where the above models are prompted to avoid obvious LLM-writing indicators.
Retrieval-Augmented QA. MS MARCO (Bajaj et al., 2016) is a large-scale machine-reading-comprehension benchmark comprising real Bing search queries. Each example contains an anonymized user query, a set of candidate web passages retrieved by the search engine, and a human-written answer. We randomly sample 100 queries and generate responses from Llama-4 Scout (AI, 2024), OLMo-2-13B Instruct (OLMo et al., 2025), Mistral-7B (Jiang et al., 2023), GPT-4o (OpenAI et al., 2024), and Gemma-2-27B (Team et al., 2024) (Generation details in Appendix C).
@@ 3.2 Pilot “Slop” Annotation
As an initial validation of the taxonomy, we hire 5 professional copy-editors from the Upwork platform [^4] https://www.upwork.com/ to annotate “slop” spans in two datasets: News article generation (Russell et al., 2025); and Retrieval Augmented QA (MS MARCO; Bajaj et al., 2016). [^5] Exempt Determination obtained from our institutional IRB. See Ethical Considerations for approval details. These datasets span different writing styles, expected passage lengths, and serve different purposes for the reader. We paid annotators at a rate of \$35-45 USD an hour. Each article took $\sim$10-15 minutes to annotate, with an average of 871 words per article. The MS MARCO dataset took $\sim$4-7 minutes to annotate, with an average of 55 words per passage. We provided annotators with a guide containing codes for indicators of “slop” from our expert definition survey (§3; Appendix Table 6). We asked annotators to read the text in full, and first answer whether they initially perceive the text as “slop.” We then had annotators label spans in the text (word-level) that instantiate any of the “slop” codes. Texts may have span annotations even if not initially deemed by the annotator to be “slop” overall. We provide the full set of questions given to annotators in Appendix Figure 5.666We built the labelling interface using a custom template in LabelStudio: https://labelstud.io/. In sum, each annotator annotated the same 10 articles and 10 passages in the pilot round.
@@ 3.3 Finalized “Slop” Taxonomy
Here, we describe each theme and code after annotator adjudication (See Appendix F for a description with examples).
Information Utility assesses how effectively a text conveys meaningful and contextually appropriate information. It comprises two key indicators: (i) Density, or the amount of substantive content relative to the length of the text, measured through information-theoretic token entropy (Meister et al., 2021) and propositional idea density (Brown et al., 2008), and (ii) Relevance, the alignment of content with task or prompt, measured through expert human annotations due to complexities in automated assessments (Clarke & Dietz, 2024).
Information Quality describes the accuracy and subjectivity of the presented information. Factuality assesses inaccuracies, hallucinations, or fallacious claims within the text, which require human annotations due to the complexity of automated factual evaluations in the absence of reference texts (Ramprasad & Wallace, 2024; Laskar et al., 2023). Bias (Subjectivity) assesses the presence or absence of a necessary subjective or rhetorical perspective, measured by the proportion of subjective words through an established lexicon (Wiebe et al., 2004).
Style Quality addresses properties related to expression and readability. Repetition, identified by lexical repetition metrics (Shaib et al., 2024a) and Templatedness, measured via syntactic structures (Shaib et al., 2024b) are key features of text Structure. Coherence is evaluated mostly via expert annotations due to the absence of reliable automatic measurements (Li et al., 2024; Murugadoss et al., 2024). Aspects of Tone evaluate the appropriateness and character of language relative to context, and include issues like excessive formality (Fanous et al., 2025; Yang et al., 2024). We include indicators such as Fluency (naturalness of language); Verbosity (passage and sentence length) (Zhang et al., 2024); and Word Complexity, i.e., use of unnecessarily complex vocabulary, measured by Gunning-Fog Index (Gunning, 1952) and Flesch-Kincaid Grade Level (Kincaid et al., 1975).
@@ 5 What is “Sloppy” Text?
We now present results from all annotations collected across the news and QA datasets. Our analysis includes both a combined setting across all domains, as well as separate evaluations by domain (news vs. QA). For each setting, we report results aggregated across all annotators. (See Appendix F for individual plots). The combined analysis highlight common slop features shared across all data and annotators, while the disaggregated evaluations show variations that may arise from annotator subjectivity or domain-specific patterns.
Using all defined “slop” features, we fit a logistic regression model [^7] Using statsmodels, version 0.14.0. with these features as the independent variables and the binarized slop label as the (single) dependent variable. Features with adjusted $p<0.05$ (after Bonferroni correction) are considered statistically significant predictors of whether annotators label texts as “slop.” Figure 2 shows that the individual axes influencing slop assessments vary slightly in the domain-specific regressions.
@@ 5.1 Results
Factuality “[…] leading to more frequent and severe heatwaves. “Climate change is like adding steroids to our weather,” says Dr. Michael Oppenheimer, a climate scientist at Princeton.
Structure “But did you know there’s another important number-sort of like a ”secret” code—printed just beneath the sell-by date? […] Find the secret code, which is usually near the sell-by date.
Tone “The very power of the word [“witch”] lies in its imprecision. It is not merely a word but an archetype, a cluster of powerful images.” The uncertainty of exactly what a witch is forms part of the titillation
@@
Disaggregated analysis shows that features important for “slop” vary based on the purpose of the text: Factual and structural issues are significant for QA data, while stylistic and utility issues are prominent for news articles. This distinction indicates the importance of evaluating LLM-written texts with respect to domain to contextualize their quality.
@@ 6.3 Can LLMs Self Identify “Slop”?
Recent text evaluations have prompted LLMs to judge text qualities not readily captured by existing metrics (e.g., Liu et al. 2023; Zheng et al. 2023). This is usually done zero-shot, providing instructions for evaluation. Here we test three LLMs (GPT-5, Deepseek-V3, and o3-mini) for their ability to (a) predict binary “slop” labels and (b) extracting “slop” spans. In both cases, we provide the full “slop” guide given to annotators (See Appendix H for the full prompt given to each model).
@@
Results: Extracting “Slop” Spans On average, GPT-5 extracted longer text spans than human annotators (mean 58 vs. 41 characters, respectively). However, span-level alignment with human annotations is low. Table 7 shows the results of prompting GPT-5 zero-shot to extract spans, and with in-context examples ($k\in[1,3,5]$). We report the character-level precision and recall. GPT-5 achieves a precision of 0.08, recall of 0.12. Providing examples up to $k=5$ does not improve precision nor recall by much: reporting precision of 0.13, and recall of 0.19. Further, the additional in-context examples do not significantly change which spans are extracted, there is a consistent F1 overlap of 0.65-0.67 between each $k$ setting and $k=0$. While the higher recall relative to precision suggests the model can identify some relevant text spans, the overall overlap remains low.

### v2

@@ Abstract
AI “slop” is an increasingly popular term used to describe low-quality AI-generated text, but there is currently no agreed upon definition of this term nor a means to measure its occurrence. In this work, we develop a taxonomy of “slop” through interviews with experts in NLP, writing, and philosophy, and propose a set of interpretable dimensions for its assessment in text. Through span-level annotation, we find that binary “slop” judgments are (somewhat) subjective, but such determinations nonetheless correlate with latent dimensions such as coherence and relevance. Our framework can be used to examine AI-generated text in both detection and binary preference tasks, potentially offering new insights into the linguistic and stylistic factors that contribute to quality judgments. We highlight that fully automated and scalable methods remain an open challenge.
@@ 1 Introduction
“Slop” has emerged as a term describing generic, low-quality content that appears to have been generated by AI. [^1] Slop was on the shortlist of Oxford Dictionary’s Word of the Year 2024, which claims a “332% increase” in usage: https://corp.oup.com/word-of-the-year/#shortlist-2024 Recent news articles offer salient examples of such AI “slop”, ranging from non-factual claims (“… add nontoxic glue to make cheese stick to a pizza”, “geologists advise eating at least one rock a day”; Hoffman, 2024; Wallace-Wells, 2024) to useless information (“fodder for websites whose only purpose appears to be optimising for [search engines]”; Mahdawi, 2025). Conversations on social media highlight indicators of “slop” in LLM responses, including overuse of certain terms, low information density, and structural quirks such as lists-as-responses. [^2] https://x.com/aidan_mclau/status/1884770586276381179 Despite the sudden ubiquity of the term, there is no clear definition of, nor method, for measuring “slop” in text.
@@
Our main contributions are as follows: We first introduce a working definition and taxonomy of “slop” and map each dimension to automatic metrics where possible (§3). To validate this framework, we collect span-level annotations from expert writers over 150 news articles and 100 question-answering passages to provide a fine-grained analysis of slop indicators (§4). Although binary assessments of “slop” vary across individuals, we show that our taxonomy helps explain which latent qualities (e.g., coherence, relevance, structural features) contribute most to these judgments5). We also find that the strength of latent qualities vary based on domain, and that our taxonomy provides a useful framework for guiding quality assessment under different tasks5). For axes that lack adequate automatic measurements (e.g., relevance, coherence, fluency), we find that standard text metrics fail to capture annotator preferences. Finally, we show that capable reasoning LLMs also fail to reliably extract and identify “slop” in text6). We provide the annotation guidelines and data at https://github.com/cshaib/slop.
@@ 2 Related Work
Text Diversity. Prior work has sought to characterize aspects of texts related to how repetitive and templated they are. Salkar et al. (2022) investigated repeated $n$-grams in LLM outputs in the context of summarization. Shaib et al. (2024b) found that modern LLMs are prone to repeatedly generate favoured syntactic templates, i.e., sequences of Part-of-Speech (PoS) tags. Padmakumar and He (2024) and Tevet and Berant (2020) examined lexical and semantic diversity in generated texts, introducing metrics to quantify variation across outputs and emphasizing its importance to generation quality. These existing efforts have informed the way in which we are thinking about what characterizes writing style and AI “slop” and provides automatic measurements for key aspects of “slop.”
Text Quality Measurements. Text quality has typically been measured using simple surface-level metrics like BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004), which can be effective when reference outputs are available (e.g., machine translation evaluation). More recent work has recognized that text quality is not monolithic but rather comprises multiple, sometimes competing dimensions that must be measured independently, and accordingly focused on multidimensional frameworks assessing properties of texts. Chakrabarty et al. (2025b) provide an editing taxonomy to correct (Chakrabarty et al., 2025a) recurring AI-writing flaws such as cliches and unnecessary exposition. Similarly, Bharadwaj et al. (2025) show that reward models over-weight 5 superficial writing cues including length, structure, jargon, sycophancy, and vagueness. Both works confirm that multiple factors contribute to text quality. Our work is complementary to measuring quality in general: We target stylistic patterns unique to LLM writing that are not covered by other taxonomies.
@@ 3 Defining LLM “Slop”
The Oxford Dictionary defines “slop” as: “[…] material produced using a large language model (LLM), which is often viewed as being low-quality or inaccurate. This type of low-quality, AI-generated material is becoming increasingly visible to people […], who often view it as unwanted or inferior.”
“Slop” as a construct does not immediately permit measurement: It is difficult to quantify “low-quality” or “unwanted” text. We propose a composite measure over observable characteristics of text, where we elicit salient characteristics from a set of individuals with a range of relevant expertise. Human writing can also read as “slop”, but we adopt the above definition and focus on (seemingly) LLM-generated texts. [^3] Interestingly, what is considered “slop” in human-written text can differ in characteristics, and may even serve as an intermediary in writing processes (cf. Appendix B)
@@
Using qualitative content analysis and deductive coding techniques, we map expert definitions of “slop” to measurable concepts (Hsieh and Shannon, 2005). We begin by identifying key terms in survey responses and building a code list until saturation (i.e., until no new codes are created). We then map each response on to one of the following codes: Factuality, Information Density, Bias, Relevance, Repetition, Templatedness, Verbosity, Word Complexity, Tone, Coherence, Fluency, Diversity, Engagement, Vagueness, and Utility.
@@ 3.1 Datasets
We select two datasets to annotate for “slop”, motivated by two practical observations: First, that LLM-written text is becoming commonplace in reporting news on the internet (Tomlinson et al., 2025; Chatterji et al., 2025a). The second is motivated by the use-cases outlined by our experts in §3, where a majority ($n=9$) reported using LLMs for question answering tasks (Appendix B).
News Article Generation. We evaluate “slop” over 150 news articles released by Russell et al. (2025), in which annotators are asked to label texts as being AI-written or not. Each unique article has a human-written source and a corresponding AI-written article, generated by either Claude (Anthropic, 2024), GPT-4o or o1-pro (OpenAI et al., 2024). Additionally, each article includes a “humanized” article, where the above models are prompted to avoid obvious LLM-writing indicators.
Retrieval-Augmented QA. MS MARCO (Bajaj et al., 2018) is a large-scale machine-reading-comprehension benchmark comprising real Bing search queries. Each example contains an anonymized user query, a set of candidate web passages retrieved by the search engine, and a human-written answer. We randomly sample 100 queries and generate responses from Llama-4 Scout (AI, 2024), OLMo-2-13B Instruct (OLMo et al., 2025), Mistral-7B (Jiang et al., 2023), GPT-4o (OpenAI et al., 2024), and Gemma-2-27B (Team et al., 2024) (Generation details in Appendix C).
@@ 3.2 Pilot “Slop” Annotation
As an initial validation of the taxonomy, we hire 5 professional copy-editors from the Upwork platform [^4] https://www.upwork.com/ to annotate “slop” spans in two datasets: News article generation (Russell et al., 2025); and Retrieval Augmented QA (MS MARCO; Bajaj et al., 2018). [^5] Exempt Determination obtained from our institutional IRB. See Ethical Considerations for approval details. These datasets span different writing styles, expected passage lengths, and serve different purposes for the reader. We paid annotators at a rate of \$35-45 USD an hour. Each article took $\sim$10-15 minutes to annotate, with an average of 871 words per article. The MS MARCO dataset took $\sim$4-7 minutes to annotate, with an average of 55 words per passage. We provided annotators with a guide containing codes for indicators of “slop” from our expert definition survey (§3; Appendix Table 7). We asked annotators to read the text in full, and first answer whether they initially perceive the text as “slop.” We then had annotators label spans in the text (word-level) that instantiate any of the “slop” codes. Texts may have span annotations even if not initially deemed by the annotator to be “slop” overall. We provide the full set of questions given to annotators in Appendix Figure 6.666We built the labelling interface using a custom template in LabelStudio: https://labelstud.io/. In sum, each annotator annotated the same 10 articles and 10 passages in the pilot round.
@@ 3.3 Finalized “Slop” Taxonomy
Here, we describe each theme and code after annotator adjudication (See Appendix I for a description with examples).
Information Utility assesses how effectively a text conveys meaningful and contextually appropriate information. It comprises two key indicators: (i) Density, or the amount of substantive content relative to the length of the text, measured through information-theoretic token entropy (Meister et al., 2021) and propositional idea density (Brown et al., 2008), and (ii) Relevance, the alignment of content with task or prompt, measured through expert human annotations due to complexities in automated assessments (Clarke and Dietz, 2024).
Information Quality describes the accuracy and subjectivity of the presented information. Factuality assesses inaccuracies, hallucinations, or fallacious claims within the text, which require human annotations due to the complexity of automated factual evaluations in the absence of reference texts (Ramprasad et al., 2024). Bias (Subjectivity) assesses the presence or absence of a necessary subjective or rhetorical perspective, measured by the proportion of subjective words through an established lexicon (Wiebe et al., 2004).
Style Quality addresses properties related to expression and readability. Repetition, identified by lexical repetition metrics (Shaib et al., 2024a) and Templatedness, measured via syntactic structures (Shaib et al., 2024b) are key features of text Structure. Coherence is evaluated mostly via expert annotations due to the absence of reliable automatic measurements (Li et al., 2024; Murugadoss et al., 2025). Aspects of Tone evaluate the appropriateness and character of language relative to context, and include issues like excessive formality (Fanous et al., 2025; Yang et al., 2024). We include indicators such as Fluency (naturalness of language); Verbosity (passage and sentence length) (Zhang et al., 2024); and Word Complexity, i.e., use of unnecessarily complex vocabulary, measured by Gunning-Fog Index (Gunning, 1952) and Flesch-Kincaid Grade Level (Kincaid et al., 1975).
@@ 5 What is “Sloppy” Text?
We now present results from all annotations collected across the news and QA datasets. Our analysis includes both a combined setting across all domains, as well as separate evaluations by domain (news vs. QA). For each setting, we report results aggregated across all annotators. (See Appendix I for individual plots). The combined analysis highlight common slop features shared across all data and annotators, while the disaggregated evaluations show variations that may arise from annotator subjectivity or domain-specific patterns.
We first construct “slop” features as an aggregated count and presence of the span-level codes across annotators. We fit a logistic regression model [^7] Using statsmodels, version 0.14.0. with these features as the independent variables and the binarized slop label as the (single) dependent variable. This allows us to evaluate whether aggregated patterns in the span-level taxonomy are associated with the binary slop judgments. Features with adjusted $p<0.05$ (after Bonferroni correction) are considered statistically significant predictors of whether annotators label texts as “slop.” Figure 2 shows that the individual axes influencing slop assessments vary slightly in the domain-specific regressions.
@@ 5.1 Results
Factuality “[…] leading to more frequent and severe heatwaves. “Climate change is like adding steroids to our weather," says Dr. Michael Oppenheimer, a climate scientist at Princeton.
Structure “But did you know there’s another important number-sort of like a "secret" code—printed just beneath the sell-by date? […] Find the secret code, which is usually near the sell-by date.
Tone “The very power of the word [“witch"] lies in its imprecision. It is not merely a word but an archetype, a cluster of powerful images." The uncertainty of exactly what a witch is forms part of the titillation
@@
Disaggregated analysis shows that features important for “slop” vary based on the purpose of the text: Factual and structural issues are significant for QA data, while stylistic and utility issues are prominent for news articles. This distinction indicates the importance of evaluating LLM-written texts with respect to domain to contextualize their quality. [^8] We provide further qualitative assessment of the label distribution across topics and sources within the news domain in Appendix H.
@@ 6.3 Can LLMs Self Identify “Slop”?
Recent text evaluations have prompted LLMs to judge text qualities not readily captured by existing metrics (e.g., Liu et al. 2023; Zhang et al. 2024). This is usually done zero-shot, providing instructions for evaluation. Here we test three LLMs (GPT-5, Deepseek-V3, and o3-mini) for their ability to (a) predict binary “slop” labels and (b) extracting “slop” spans. In both cases, we provide the full “slop” guide given to annotators (See Appendix L for the full prompt given to each model).
@@
6.3.1 LLM-Generated Span Rationales
In addition to extracting spans, the LLM-as-a-judge is instructed to provide a rationale for each selection. Reasoning chains do not reliably return the exact codes from the guide, so to assess the overall alignment of the reasoning and the human-assigned codes, we count and rank tri-grams from the judge-generated rationale (Figure 3, Appendix Figure 8).
Figures 3(a) and 3(b) highlight a mismatch between the human-assigned codes and those flagged most frequently by o3-mini. Specifically, o3-mini overwhelmingly focuses on Density-related issues, and does not show the full range of codes used by annotators. In Appendix G, we show that this does not significantly change when in-context examples are provided to the model.
Figure 3: Frequency of (a) human-assigned slop code prevalence and, (b) “slop” category collapsed trigrams in o3-mini span rationales in the News domain. o3-mini over-represents issues related to Density in the rationales, while under-representing issues with Coherence, Tone, Bias, Relevance, and Factuality.
Results: Extracting “Slop” Spans On average, GPT-5 extracted longer text spans than human annotators (mean 58 vs. 41 characters, respectively). However, span-level alignment with human annotations is low. Table 7 shows the results of prompting GPT-5 zero-shot to extract spans, and with in-context examples ($k\in[1,3,5]$). We report the character-level precision and recall. [^9] Implementation details in Appendix K. GPT-5 achieves a precision of 0.08, recall of 0.12. Providing examples up to $k=5$ does not improve precision nor recall by much: reporting precision of 0.13, and recall of 0.19. Further, the additional in-context examples do not significantly change which spans are extracted, there is a consistent F1 overlap of 0.65-0.67 between each $k$ setting and $k=0$. While the higher recall relative to precision suggests the model can identify some relevant text spans, the overall overlap remains low.